# TWIN-GPT: Digital Twins for Clinical Trials via Large Language Model

### v1

@@ Abstract
Clinical trials are indispensable for medical research and the development of treatments (e.g., drugs). However, a clinical trial often necessitates the involvement of hundreds of participants and can span several years to complete, with a high probability of failure during the process. Recently, there has been a burgeoning interest in virtual clinical trials, which simulate real-world scenarios and hold the potential to significantly enhance patient safety, expedite development, reduce costs, and contribute to the broader scientific knowledge in healthcare. Existing research often focuses on leveraging electronic health records (EHRs) to support clinical trial outcome prediction. Yet, trained with limited clinical trial outcome data, existing approaches frequently struggle to perform accurate predictions. Some research has attempted to generate EHRs to augment model development, but has fallen short in personalizing the generation for individual patient profiles. Recently, the emergence of large language models has illuminated new possibilities, as their embedded comprehensive clinical knowledge has proven beneficial in addressing medical issues. In this paper, we propose a large language model-based digital twin creation approach, called TWIN-GPT. TWIN-GPT can establish cross-dataset associations of medical information given limited data, generating unique personalized digital twins for different patients, thereby preserving individual patient characteristics. Comprehensive experiments show that, using digital twins created by TWIN-GPT can boost the clinical trial outcome prediction, exceeding various previous prediction approaches. Besides, we also demonstrate that TWIN-GPT can generate high-fidelity trial data that closely approximates specific patients, aiding in more accurate result predictions in data-scarce situations. Moreover, our study provides practical evidence for the application of digital twins in healthcare, highlighting its potential significance.
@@ 1 Introduction
Clinical trials serve as essential research investigations in the field of prospective medicine, designed to evaluate the safety and efficacy of new treatments (e.g., drugs, or other interventions) rigorously. Despite playing a pivotal role in advancing medical knowledge and enhancing patient care, these meticulously orchestrated studies typically engage tens to thousands of human participants and frequently extend over several years, hindering the progress of research and development. There is an increasing interest in leveraging artificial intelligence models to predict trial outcomes, leveraging electronic health records (EHR) and significantly accelerating the pace of research [21, 56, 23, 33, 19]. Learning from EHR data, models are able to identify the relationship between potential health issues with the personalized characteristics, thus excelling at forecasting disease occurrence and even developing personalized treatment plans [59, 47].
@@
Current approaches leveraging electronic health records (EHR) for clinical trial outcome prediction still face a set of challenges. These methods often fail to account for individual differences between patients and variations among diseases, resulting in predictions that may not align with real-world scenarios [33, 19]. Current EHR-based clinical trial outcome prediction models typically suffers from two key problems: (1) Data Gap: the disconnect between the data and its related realistic background knowledge makes it hard to leverage the supportive knowledge in prediction. (2) Data Inconsistency: discrepancies between data from different sources, presenting different data distributions, different patterns of data missing, and different data recording formats, undermine the reliability of prediction results across datasets.
@@
Recently, digital twin creation technology has exhibited great potential in clinical practice [15, 26], simulating patient physiological state [41, 11], disease progression [76], and treatment effects [1, 13]. These technologies leverages historical and real-time data to provide medical professionals with more accurate and comprehensive personalized health information, introducing opportunities in outcome prediction and analysis. Inspired by this, Das et al[27] designed a simple model for cancer patients’ EHR digital twin creation, which promotes privacy protection and the outcome prediction in clinical trial simulation. However, this method only relies on the existing datasets when creating digital twins, which fails to leverage knowledge beyond the dataset and limits the accuracy of clinical trial outcome prediction. Traditional digital twin models still face challenges in achieving accurate clinical trial outcome predictions, failing to effectively address data gaps and data inconsistencies. Data gaps are particularly prominent in trial predictions due to variations in ICD codes for different medications and diseases, necessitating the acquisition of real-world background knowledge by the digital twin for clinical trials. Moreover, the data inconsistency problem is greatly exacerbated in EHR data for clinical trial scenarios, posing a challenge to digital twin creation, since EHRs for clinical trials often originate from different sites and thus present greater inconsistencies.
@@ 2.1 Large Language Models (LLMs) for Medicine
Recently, LLMs have exhibited excellent performance on various tasks, especially in the field of medicine [62, 25, 45]. The current research on LLMs in the medical field has provided novel applications in bedside diagnosis, automated drafting of clinical documents, and improving medical workflow efficiency [12, 54, 51]. LLMs have demonstrated their utility in clinical applications by passing medical licensing exams and offering superior empathy and quality of medical advice compared to human clinicians [69, 42, 52, 61, 71, 75, 73]. In clinical applications, LLMs were used to facilitate the analysis of large amounts of textual data, providing insights that can drive medical discoveries and increase the efficiency of research workflows [66, 67, 60, 4]. They also have significant advantages in daily clinical tasks like clinical documentation drafting, and can significantly reduce the administrative burden on medical staff [78, 36, 10, 30]. However, there is no research using LLMs in clinical trial scenarios, which is a crucial area demanding the extensive medical knowledge encoded within these large language models.
@@ 2.2 Patient Outcome Prediction
Previous research in patient outcome prediction primarily relied on traditional machine learning methods such as decision trees, support vector machines, and logistic regression [20, 40, 14]. These approaches often utilized manually selected features from high-dimensional clinical data. However, manual feature selection frequently fails to incorporate informative features, thereby failing to fully capture individualized patient information and complex data relationships [22]. Recently, deep learning methods have gained prominence with the advancement of deep learning technology. Researchers have begun to explore the use of neural networks for patient prediction [21, 56, 23, 44], which handle well with large-scale data and automatically select informative features, leading to significant achievements. Fu et al. [33, 32] proposed a hierarchical interaction network (HINT) to predict the clinical trial outcome based on drug molecules, disease codes, eligibility criteria, etc. Lu et al. [48] enhance HINT from the perspective of uncertainty quantification and interpretability.
Among these, analysis of EHR data plays a crucial role in patient outcome prediction. Many studies have investigated how to effectively utilize EHR data for patient prediction [20, 39, 55], applying long-term historical EHR learning, medical text mining, time-series data analysis [43, 2, 79]. Utilizing EHR data can provide comprehensive patient information, but it also suffer from data gap problem and data inconsistency problem. Recently, digital twin creation has emerged as a noteworthy field in EHR learning, offering new opportunity to patient outcome prediction [63]. Digital twin models can simulate a patient’s personalized physiological states, disease progression, and treatment effects based on historical EHR and real-time data [27].
@@ 2.3 EHR Data Generation
Digital twin is a specialized method for generating personalized EHR data, ensuring similarity and correspondence with the original data, as well as guaranteeing patient privacy protection [49]. Das et al[27] used a VAE to model digital twins in clinical trials to achieve prediction of patient outcomes in low-data scenarios. Liu et al[46] proposed a cloud healthcare system framework based on digital twin healthcare (CloudDTH), which is used to monitor, diagnose, and predict personal health to achieve the goal of personal health management. Zhong et al[81] used ICU EHR data to build the ICU digital twin model to study critical care services in the ICU. However, these approaches only perform digital twinning within their datasets and do not solve the data gap and data inconsistency problems.
@@ 3 METHODOLOGY
In this section, we introduce TWIN-GPT, a novel approach that creates personalized digital twins for effective virtual clinical trial simulation, leveraging the power of large language models (LLMs) to address the challenges of clinical trial outcome predictions in limited data scenarios. Our method is predicated on the understanding that clinical trials are critical for medical research and drug development but are hampered by the extensive time and participant involvement required. TWIN-GPT aims to revolutionize this by generating high-fidelity, personalized digital twins of patients, thus facilitating virtual clinical trials that can significantly enhance patient safety, expedite development, and reduce costs.
In Section 3.1, we define the problem of creating a digital twin model for clinical trials. Section 3.2 details the intricacies of data input encoding and the data preparation necessary for effective model training. In Section 3.3, we introduce training strategies for synthesizing digital twins from patient data. Then, Section 3.4 elaborates on the training process for the digital twin model, enhancing its adaptability to individual patient data. Finally, Section 3.5 discusses the generation of digital twins, illustrating how the model manipulates synthesized data to simulate realistic EHR data.
3.1 Problem Formulation
@@
We assume that there are a total of $N$ participants in a clinical trial. Each participant $n$’s clinical encounters are encapsulated as
$X_{n;1:T_{n}}=\left\{x_{n,1},x_{n,2},\dots,x_{n,T_{n}}\right\},$ (1)
@@
where each $x_{n,t}$is a collection of event types occurring at the $t$-th visit, further delineated as
$x_{n,t}=\left\{x_{n,t}^{1},x_{n,t}^{2},\dots,x_{n,t}^{U}\right\},$ (2)
$x_{n,t}$ capturing the diverse clinical events ranging from treatments to adverse reactions. where $U$ is the total number of all types of events. An event set $x_{n,t}^{u}$ contains all u type events that occurred at the $t$-th visit for the patient $n$. $x_{n,t}^{u}=\left\{c_{1},c_{2},\dots,c_{n}\right\}$ represents all events of type u, and the value of $c_{l}$ is 1 or 0, indicating whether the events happened or not. Utilizing TWIN-GPT, we synthesize digital twins $\hat{X}_{n;1:T_{n}}$ by generating personalized patient trajectories based on their real-world counterparts and nearest neighbor patients, encapsulating a wide spectrum of medical conditions and treatments. Formally, this is represented as:
$\hat{X}_{n;1:T_{n}}=\text{TWIN-GPT}\left(X_{n;1:T_{n}},\{X_{n^{\prime};1:T_{n^{\prime}}}\}_{n^{\prime}=1}^{N^{\prime}},\Theta\right),$ (3)
where $\Theta$ denotes the learned model parameters, and $N^{\prime}$ is the count of nearest neighbor patients considered for synthesizing the digital twin. Then we identify the most similar feature vector from the dataset, which will serve as the ground truth:
$GT_{n,t}={sim}\left(\bar{F}_{n,t},\{F_{n^{\prime};t^{\prime}}\}\right),$ (4)
where $GT_{n,t}$ represents the ground truth for patient $n$ at visit $t$, derived from the pool of averaged feature vectors $\{F_{n^{\prime};t^{\prime}}\}$ across all patients and visits.
In our analysis, we particularly focus on three critical aspects: treatment approaches, medication usage, and adverse events. Treatment is the assigned treatment at the $n$-th timestep for the participant $n$, denoted $x_{n,t}^{\text{treat}}$. medication denoted as $x_{n,t}^{\text{med}}$, is defined as the treatment drugs given to patients in response to their current health conditions. Adverse events represented by $x_{n,t}^{\text{ae}}$, refer to all unexpected health incidents that occur during patient visits. To let the model predict the adverse events in the next visit as $X_{n,t+1}^{\text{ae}}$, we choose to input all three vital type events represented by $X_{n,t}^{\text{treat}}$, $X_{n,t}^{\text{med}}$ and $X_{n,t}^{\text{ae}}$ because the causality between three types of events play an important role in predicting. For example, the doctor might provide medication Docetaxel to the patient at timestep $t+1$ because of the adverse events observed in the last visit. Meanwhile, the adverse event pain might not be observed at timestep $t+1$ due to the medication Docetaxel given at timestep $t$. It is easy for large language models, such as ChatGPT, to fully capture the causality between all kinds of events because they have robust in-context learning ability [50].
3.2 Input Encoding
In the process of clinical trials, we define the original patient record $X_{n;1:T_{n}}$ by collecting the sequence of patient visit records. Our goal is to generate a digital twin $\hat{x}_{n;1:T_{n}}$ that can retain the features of the target patient. For this purpose, we propose a generator, namely TWIN-GPT, that simulates personalized patient trajectories, which are determined by the real record $X_{n;1:T_{n}}$, $\{X_{k;1:T_{n}}\}_{k}$ is the nearest $K$ neighbor patients, and $\theta$ is the learned model parameters that are constantly updated during training, as shown below:
$\hat{X}_{n;1:T_{n}}=\textit{TWIN-GPT}(X_{n;1:T_{n}},\{X_{k;1:T_{k}}\}_{k};\theta).$ (5)
Each visit $x_{n,t}$ within the primary input $X_{n;1:T_{n}}$ is structured from $U$ diverse concurrent event types, illustrated as $x_{n,t}=\{x^{1}_{n,t},\ldots,x^{U}_{n,t}\}$, forming a chronological sequence of visits $x_{n,t}$.The event set $x^{u}_{n,t}$ is converted into a multi-hot vector, where each element represents the occurrence of a specific event at this time step, i.e., $x^{u}_{n,t}=\{c_{1},\ldots,c_{l}\}$, where $c_{i}\in\{0,1\}(i=1,2,3,...,l)$.
The initial input $X_{n;1:T_{n}}$ is constructed from a series of patient visits $x_{n,t}$, each visits $x_{n,t}$ encompasses $U$ distinct categories of events simultaneously, such as $x_{n,t}=\{x^{1}_{n,t},\ldots,x^{U}_{n,t}\}$. The event set $x^{u}_{n,t}=\{c_{1},\ldots,c_{l}\}$ is converted into a multi-hot vector, where each element represents the occurrence of a specific event at this time step, where $c_{i}\in\{0,1\}(i=1,2,3,...,l)$.
3.3 Prompt Tasks
The prompt task is a piece of text input provided to the model to generate a response based on its training. We proposed a unique approach to train Electronic Health Record (EHR) model. This approach is based on individual patient sessions and generates a digital twin of the patient through a series of steps. The process includes the following stages:
(1) Initially, a digital twin is established based on the patient’s EHR data. At this stage, the patient’s EHR data and five EHR datasets similar to the patient’s condition are provided. These data are input into the process of generating the digital twin. (2) Next, the correct answer is provided, with the expectation that the training process will be optimized by summarizing and generalizing these answers. This process is iterated according to the number of patient visits. For each visit, the patient’s profile and the EHR data from the first visit are provided, followed by five EHR profiles similar to the patient’s symptoms. Based on this, an adverse event that the patient may experience under the current treatment is predicted. If the predicted outcome deviates significantly from the correct answer, multiple-choice answers will be provided [37]. (3) Subsequently, the correct adverse event under the current treatment is provided. Based on these adverse events and the treatment, the medication that the patient should use next time is deduced and this information is provided. (4) Finally, a digital twin of the patient’s medication and adverse events is generated. This digital twin must take into account the specific treatment name.

### v2

@@ Abstract
Clinical trials are indispensable for medical research and the development of new treatments. However, clinical trials often involve thousands of participants and can span several years to complete, with a high probability of failure during the process. Recently, there has been a burgeoning interest in virtual clinical trials, which simulate real-world scenarios and hold the potential to significantly enhance patient safety, expedite development, reduce costs, and contribute to the broader scientific knowledge in healthcare. Existing research often focuses on leveraging electronic health records (EHRs) to support clinical trial outcome prediction. Yet, trained with limited clinical trial outcome data, existing approaches frequently struggle to perform accurate predictions. Some research has attempted to generate EHRs to augment model development but has fallen short in personalizing the generation for individual patient profiles. Recently, the emergence of large language models has illuminated new possibilities, as their embedded comprehensive clinical knowledge has proven beneficial in addressing medical issues. In this paper, we propose a large language model-based digital twin creation approach, called TWIN-GPT. TWIN-GPT can establish cross-dataset associations of medical information given limited data, generating unique personalized digital twins for different patients, thereby preserving individual patient characteristics. Comprehensive experiments show that using digital twins created by TWIN-GPT can boost the clinical trial outcome prediction, exceeding various previous prediction approaches. Besides, we also demonstrate that TWIN-GPT can generate high-fidelity trial data that closely approximates specific patients, aiding in more accurate result predictions in data-scarce situations. Moreover, our study provides practical evidence for the application of digital twins in healthcare, highlighting its potential significance.
@@ 1 Introduction
Clinical trials serve as essential research investigations in the field of prospective medicine. They are designed to rigorously evaluate the safety and efficacy of new treatments (e.g., drugs, or medical devices). Despite playing a pivotal role in advancing medical knowledge and enhancing patient care, these trials typically engage tens to thousands of human participants and frequently extend over several years, hindering the progress of research and development. There is an increasing interest in leveraging artificial intelligence models to predict trial outcomes, leveraging electronic health records (EHR) and significantly accelerating the pace of research [58, 26, 36, 22]. Learning from EHR data, models are able to identify the relationship between potential health issues with personalized characteristics, thus excelling at forecasting disease occurrence and even developing personalized treatment plans [61, 49].
@@
Existing approaches leveraging EHR for clinical trial outcome prediction still face a set of challenges. These methods often fail to account for individual differences between patients and variations among diseases, resulting in predictions that may not align with real-world scenarios [36, 22]. EHR-based clinical trial outcome prediction models typically suffer from two key problems: (1) Data Gap: the disconnect between the data and its related realistic background knowledge makes it hard to leverage the supportive knowledge in prediction. (2) Data Inconsistency: discrepancies between data from different sources, presenting different data distributions, different patterns of data missing, and different data recording formats, undermine the reliability of prediction results across datasets.
@@
Digital twin creation technology has exhibited great potential in clinical practice [15, 29], simulating patient physiological state [43, 11], disease progression [78], and treatment effects [1, 13]. These technologies leverage historical and real-time data to provide medical professionals with more accurate and comprehensive personalized health information, introducing opportunities in outcome prediction and analysis. Inspired by this, Das et al[30] designed a simple model for cancer patients’ EHR digital twin creation, which promotes privacy protection and the outcome prediction in clinical trial simulation. However, this method only relies on the existing datasets when creating digital twins, which fails to leverage knowledge beyond the dataset and limits the accuracy of clinical trial outcome prediction. Traditional digital twin models still face challenges in achieving accurate clinical trial outcome predictions, failing to effectively address data gaps and data inconsistencies. Data gaps are particularly prominent in trial predictions due to variations in ICD codes for different medications and diseases, necessitating the acquisition of real-world background knowledge by the digital twin for clinical trials. Moreover, the data inconsistency problem is greatly exacerbated in EHR data for clinical trial scenarios, posing a challenge to digital twin creation, since EHRs for clinical trials often originate from different sites and thus present greater inconsistencies.
@@ 2.1 Large Language Models (LLMs) for Medicine
LLMs have exhibited excellent performance on various tasks, especially in the field of medicine [64, 28, 47]. The current research on LLMs in the medical field has provided novel applications in bedside diagnosis, automated drafting of clinical documents, and improving medical workflow efficiency [12, 56, 53]. LLMs have demonstrated their utility in clinical applications by passing medical licensing exams and offering superior empathy and quality of medical advice compared to human clinicians [71, 44, 54, 63, 73, 77, 75]. In clinical applications, LLMs were used to facilitate the analysis of large amounts of textual data, providing insights that can drive medical discoveries and increase the efficiency of research workflows [68, 69, 62, 4]. They also have significant advantages in daily clinical tasks like clinical documentation drafting, and can significantly reduce the administrative burden on medical staff [80, 39, 10, 33]. However, there is no research using LLMs in clinical trial scenarios, which is a crucial area demanding the extensive medical knowledge encoded within these large language models.
@@ 2.2 Patient Outcome Prediction
Previous research in patient outcome prediction primarily relied on traditional machine learning methods such as decision trees, support vector machines, and logistic regression [23, 42, 14, 20]. These approaches often utilized manually selected features from high-dimensional clinical data. However, manual feature selection frequently fails to incorporate informative features, thereby failing to fully capture individualized patient information and complex data relationships [25]. Recently, deep learning methods have gained prominence with the advancement of deep learning technology. Researchers have begun to explore the use of neural networks for patient prediction [24, 58, 26, 46, 17, 19], which handle well with large-scale data and automatically select informative features, leading to significant achievements. Fu et al. [36, 35] proposed a hierarchical interaction network (HINT) to predict the clinical trial outcome based on drug molecules, disease codes, eligibility criteria, etc. Lu et al. [50] enhance HINT from the perspective of uncertainty quantification and interpretability.
Among these, analysis of EHR data plays a crucial role in patient outcome prediction. Many studies have investigated how to effectively utilize EHR data for patient prediction [23, 41, 57, 75], applying long-term historical EHR learning, medical text mining, time-series data analysis [45, 2, 81]. Utilizing EHR data can provide comprehensive patient information, but it also suffer from data gap problem and data inconsistency problem. Recently, digital twin creation has emerged as a noteworthy field in EHR learning, offering new opportunity to patient outcome prediction [65]. Digital twin models can simulate a patient’s personalized physiological states, disease progression, and treatment effects based on historical EHR and real-time data [30].
@@ 2.3 EHR Data Generation
Digital twin is a specialized method for generating personalized EHR data, ensuring similarity and correspondence with the original data, as well as guaranteeing patient privacy protection [51]. Das et al[30] used a VAE to model digital twins in clinical trials to achieve prediction of patient outcome in low-data scenarios. Liu et al[48] proposed a cloud healthcare system framework based on digital twin healthcare (CloudDTH), which is used to monitor, diagnose, and predict personal health to achieve the goal of personal health management. Zhong et al[84] used ICU EHR data to build the ICU digital twin model to study critical care services in the ICU. However, these approaches only perform digital twinning within their datasets and do not solve the data gap and data inconsistency problems.
@@ 3 METHODOLOGY
In this section, we introduce TWIN-GPT, a novel approach that creates personalized digital twins for effective virtual clinical trial simulation, leveraging the power of large language models (LLMs) to address the challenges of clinical trials in limited data scenarios. Our method originates from the understanding that clinical trials are essential for medical research and drug development, but they are often hindered by the extensive time and participant involvement required. TWIN-GPT aims to revolutionize this by generating high-fidelity, personalized digital twins of patients, thus facilitating virtual clinical trials that can significantly enhance patient safety, expedite development, and reduce costs.
Overview. In Section 3.1, we define the problem of digital twin creation for clinical trials. Section 3.2, We employ a novel prompt-tuning approach to generate and update patient digital twins based on their EHR data, enhancing the model’s predictive accuracy and understanding of patient outcomes. In Section 3.3, We propose TWIN-GPT to generate digital twins of patients in clinical trials, enhancing prediction accuracy and understanding of patient outcomes by leveraging historical and similar patient data.
3.1 Problem Formulation and Overview Pipeline
@@
We assume that there are a total of $N$ participant patients in a clinical trial, and each patient $n$’s trial encounters are encapsulated as
$X_{n;1:T_{n}}=\left\{x_{n,1},x_{n,2},\cdots,x_{n,T_{n}}\right\},$ (1)
@@
where each $x_{n,t}$ is a collection of event types occurring at the $t$-th visit, defined as
$x_{n,t}=\left\{x_{n,t}^{1},x_{n,t}^{2},\cdots,x_{n,t}^{u}\right\}.$ (2)
Each element $x^{i}_{n,t}$ corresponds to a kind of clinical event like a treatment or an adverse reaction, and $u$ is the total number of all types of events, while $x_{n,t}^{i}=\left\{c_{1},c_{2},\cdots,c_{n}\right\}$ represents the occurrence of event, where $c_{l}$ is 1 or 0. Formally, we synthesize the digital twin of a patient by an auto-regressive model $M$, and each step predicts the next visit event:
$\hat{X}_{n,t}=M\left(X_{n^{(^{\prime})};1:t-1},\{X_{n^{\prime};1:t-1}\}_{n^{\prime}=1}^{N^{\prime}},\Theta\right),$ (3)
where $t=1,\ldots,T_{n}$, and $\Theta$ denotes the model parameters. The term $N^{\prime}$ is the count of nearest neighbor patients for reference in the digital twin generation. In the training process, we supervise the model $M$’s learning by comparing the generated digital twins with the true patient’s trial encounters during the trial. The learning objective is to maximize the cosine similarity between real and synthetic patient representations,
$\underset{\Theta}{\arg\max}\ \frac{1}{T_{n}}\sum_{t=1}^{T_{n}}{sim}(F_{n,t},\hat{F}_{n,t}),$ (4)
where ${F}_{n,t}$ and $\hat{F}_{n,t}$ denote the average representations of $x_{n,t}^{u}$ and $\hat{x}_{n,t}^{u}$ (where $\hat{x}_{n,t}^{u}$ is in $\hat{X}_{n;1}$) over the time span from $1$ to $T$. We use cosine similarity as the similarity metric (i.e., “sim” in Eq. 4).
In our analysis, we particularly focus on three essential clinical events (i.e., $u$):
  • treatment, $x_{n,t}^{\text{treat}}$, is the assigned treatment at the $t$-th timestep for the patient $n$;
  • medication, $x_{n,t}^{\text{med}}$, is defined as the treatment (i.e., drugs) given to patients in response to their current health conditions;
  • adverse event, $x_{n,t}^{\text{ae}}$, refers to all unexpected health incidents that occur during patient visits.
To predict adverse events in the next visit, denoted as $X_{n,t+1}^{\text{ae}}$, we feed all three event types into the model as the causality among these event types is crucial for accurate prediction. For example, the doctor might provide medication Docetaxel to a patient at the timestep $t+1$ due to the adverse events observed in the last visit. Meanwhile, the adverse event pain might not be observed at timestep $t+1$ due to the medication Docetaxel given at timestep $t$. Since our model is a kind of large language model, the input three event types serve as in-context learning ability [52] in our case.
3.2 Prompt Tuning for Digital Twin in LLMs
In this work, we develop a large language model (LLM) to serve as the auto-regressive model $M$ to perform the digital twin creation. To LLMs, a prompt is the initial input or query given to the model for generating a better response [83]. A context-suited prompt is beneficial to our digital twin creation task. Compared with expensive model tuning, prompt tuning is cheaper. We proposed a novel prompt-tuning approach to seek suitable prompts to generate better digital twins. The steps of our proposed prompt tuning are as follows:
  • Initially, the first version of digital twin is generated based on the patient’s EHR data. At this stage, the patient’s EHR data of previous visits and the five most similar EHR data are provided as reference.
  • Next, the real patient records (e.g., presenting the adverse event) are regarded as the ground truth to provide supervision. We use initial prompts for predicting adverse events and medications. Based on the ground truth data, the LLM’s parameters are updated. The TWIN-GPT is developed on the ChatGPT, and we utilize ChatGPT’s fine-tuning API to update it.
  • If the correct adverse event under the current treatment is predicted, the medication for the next time is deduced sequentially.